5. We have two independent samples of n observations X1,Xy,.. . ,x, and Y. Ya, . … We want to test the hypothesis Ho : μ,-μυ versus the alternative Hi : μ, μν. (a) First, assume that the null hypothesis Ho is true and find the MLE for μ (b) Then plug this estimate into the log likelihood along with the MLE μ.. and 1,-j) to calculate the LRT statistic. (e) Is this likelihood ratio test equivalent to the test that rejects the null hypothesis when (t -g)2 > e? Justify your answer.

Thank you so much!
Answer #1

7t (-10 ΣΧί-μ)2-14.-(y,-μ)2 (a) The likelihood function. L(A)--10-w20 A- n/2 exp log L(11)-1(μ)--log 140-n log π-- dl(11) 1 n7t 7t Likelihood ratio = A =-μ = exp -10 14 (cesp(勐.ifF-2 ur).m(_剞-») x-y)- ) = exp (c) We reject Ho if A<k or exp where c is obtained from size condition (-24(z-j)2) < k or, (i-j)2 > c